OpenStudy (anonymous):

pllleeeeaase help!! Solve the equation: X + 7/8=4 A. x = 4 7/8 B. x = 3 1/8 C. x = -3 1/8 D. x = -4 7/8

OpenStudy (anonymous):

B. Since you need to keep both sides of the equation equal, you need to subtract 7/8 from both sides. So x+7/8=4 => x+7/8-7/8=4-7/8 => x=3 1/8